# Topological and statistical signatures for quantifying ambiguity in semantic search

### v1

@@ 3.4 Understanding Wasserstein Distance in Topological Data Analysis
The Wasserstein distance, also known as the earth mover’s distance, quantifies the "work" required to transform one distribution into another. In topological data analysis, this measure is particularly useful when applied to persistence diagrams. The diagonal of a persistence diagram ideally represents features (often considered as noise) that appear and disappear at the same scale, thus serving as a baseline for comparison.
@@
where $N$ is the number of points in $H_{0}$, and $\gamma_{\perp}(y)$ is the y-coordinate of the orthogonal projection of each point onto the diagram’s diagonal. This calculation measures how far each point deviates from the baseline (diagonal), it implicitly weighs the features by their granularity. Each distance moved in the calculation represents a weighted adjustment of the data’s granularity, reflecting varying scales of data clustering and feature resolution. Please see the appendix for a detailed derivation of the Wasserstein distance using the Wasserstein norm for the $H_{0}$ distribution, which underscores the importance of granularity in interpreting topological data.
@@ 4 Results
In our experiments, the set $\mathcal{C}_{3}$ is initially selected as the corpus $\mathcal{C}$, with chunks from $\mathcal{C}_{10}$ used as queries set $\mathcal{Q}$. This configuration allows for the inference of query-related topics, as at least one chunk in $\mathcal{C}_{3}$ contains token sequences that are also present in the query chunks, thereby satisfying the first condition of our ambiguity definition. By design, each corpus chunk encapsulates a single topic, while query chunks can encompass one or multiple topics. We apply topic clustering to discern and isolate queries that retrieve single-topic chunks, directing our focus to those that demonstrate the ambiguity characteristic under scrutiny.
The experimental setup is reversed by selecting $\mathcal{C}_{10}$ as the corpus and $\mathcal{C}_{5}$ as the queries. In this scenario, the queries are generally associated with a single topic, though the potential for the embedding model’s semantic bias to link a small chunk query to multiple topics cannot be dismissed. Here too, topic clustering assists in filtering out queries that correlate to multiple topics, ensuring a focused analysis on relevant data points.
@@
To maintain the integrity of our experiments, we stipulate that any retrieved corpus chunk, denoted as $c_{i}=[w^{i}_{1},\ldots,w^{i}_{m}]$, must fully contain or be contained by the query chunk in both experimental configurations $(\mathcal{Q};\mathcal{C})$: $(\mathcal{C}_{10};\mathcal{C}_{3})$ and $(\mathcal{C}_{5};\mathcal{C}_{10})$. This ensures that the semantic coherence between the chunks is adequately maintained, allowing for accurate assessment of query ambiguity and topic distribution within the chunks.
@@
Figure 2: a) $W_{1}(H_{0})$ distributions for pairs of queries-corpus $(\mathcal{Q};\mathcal{C})$ chunk sets ($\mathcal{C}_{5}$;$\mathcal{C}_{10}$) and (C10;$\mathcal{C}_{3}$) in blue and red respectively for neighborhood scale $\varepsilon=0.3$. b) Max $H_{1}$ loop lifetime distributions for ($\mathcal{C}_{5}$;$\mathcal{C}_{10}$) and (C10;$\mathcal{C}_{3}$) in blue and red respectively with neighborhood scale $\varepsilon=0.3$. c)-d) $W_{1}(H_{0})$ Distribution breakdown per embedding model for (C5;C10) and (C10;C3) respectively as KDE plots. e)-f) $LT_{max}(H_{1})$ Distribution breakdown per embedding model for (C5; C10) and (C10;C3) respectively as KDE plots in log scale.
@@
Figure 3: a) Mean $W_{1}(H_{0})$ (plain lines) and 25-75% inter-quantile range (dashed lines) for ($\mathcal{C}_{5}$;$\mathcal{C}_{10}$) and ($\mathcal{C}_{10}$; $\mathcal{C}_{3}$) in blue and red respectively as a function of the relative neighborhood scaling factor $\varepsilon$. b) equivalent plot for $LT_{max}(H_{1})$.
In the experimental pairing of $\mathcal{Q}=\mathcal{C}_{5}$ and $\mathcal{C}=\mathcal{C}_{10}$, where the predefined conditions of ambiguity are not met, the $H_{0}$ signal primarily cluster around a narrow single peak at $W_{1}(H_{0})=0.6$. The absence of significant variation in neighbor arrangements, even with adjustments to the neighborhood scaling factor $\varepsilon$ (as shown in figure 3-a), indicates a lack of structured topological features such as pinched manifolds within these queries.
@@
In contrast, the pair $\mathcal{Q}=\mathcal{C}_{10}$ and $\mathcal{C}=\mathcal{C}_{3}$ demonstrates a significantly broader distribution with the mean shifted to a higher value of $W_{1}(H_{0})\approx 0.7$. This broader spread is a partial indication of a more structured neighborhood. $H_{1}$ which capture the lifetime of loops existing in the data points cloud exhibits more persistent elements as well as shown in figure 2-b. As the neighbourhood evolves from local to global (by varying $\varepsilon$), the unambiguous queries also start to exhibit persistent loops. Those loops can be interpreted as discontinuities that mark separation between different semantic clusters, reinforcing the picture of a pinched manifold that reflects a complex and discontinuous local embedding space surrounding the query chunks. This behavior is observed across different embedding models both for $H_{0}$ and $H_{1}$ demonstrating that topological signatures of ambiguity are invariant with embedding model training corpus
Figure 4: Samples joint distribution of $H_{0}$ and $H_{1}$ signatures for both unambiguous and ambiguous conditions (blue and red distributions respectively)
Notably, the ambiguous $W_{1}(H_{0})$ distribution progressively narrows and get closer to the unambiguous and as the neighborhood scaling ratio $\varepsilon$ increases, the proximity of more topics near the queries leads to an Wasserstein norm that eventually stabilizes to a point beyond which it becomes more challenging to distinguish between structured neighbourhood containing multiple topics and single topic one. Increasing the neighborhood size by tuning $\varepsilon$ also leads to more persistent loops in average for both ambiguous and unambiguous queries but with an increased scaling in the mean lifetime of ambiguous queries loops.
Finally the sample joint distribution of $W_{1}(H_{0})$ and $LT_{max}(H_{1})$ (figure 4) emphasizes the existence of long lived loops across the range covered by the homology 0 signature. Ambiguous queries contain proportionally more loops with longer lifetimes, helping to disentangle their signal from clear queries one. We can compute the empirical conditional probability density:
$p(LT_{max}(H_{1})|W_{1}(H_{0});\varepsilon)=\frac{p(W_{1}(H_{0}),LT_{max}(H_{1});\varepsilon)}{p(W_{1}(H_{0});\varepsilon)}$ (4)
Figure 5: a) Empirical conditional probability density function map $p(LT_{max}(H_{1})|W_{1}(H_{0});\varepsilon=0.3)$ for ambiguous queries only. b) Corresponding cumulative probability density function $p(LT_{max}(H_{1})>0|W_{1}(H_{0});\varepsilon=0.3)$ and $p(LT_{max}(H_{1})>0|W_{1}(H_{0});\varepsilon=4)$ as a function of $W_{1}(H_{0})$.
@@
5 Discussion and perspectives
5.1 Homology encodes ambiguity
Our findings underscore $H_{0}$ and $H_{1}$ homology signatures captured by $W_{1}(H_{0})$ and $LT_{max}(H_{1})$ as powerful tools to differentiate the two sets of query-corpus we explored. The design of our experiment aligns with our formal definition of sentence ambiguity/clarity therefore we conclude that homology capture ambiguity akin it can be used for words polysemy. The apparent similar behaviour of $H_{0}$ and $H_{1}$ features distributions across embedding models reinforces their alignment with the formal definition of ambiguity. Indeed in our definition, the notion of semantic domain or topic is an abstract concept that inevitably vary from one embedding model to another and that is approximated by clustering. Once the data filtered by neighbouring clusters, the homologies differences appear in all tested models, therefore demonstrating the general nature of homologies signature of ambiguity.
In that context, it is important to note that a finer analysis of $H_{0}$ and $H_{1}$ distribution open the door to a richer set of features for detecting and quantifying ambiguity. An example is counting the number of persistent loops per sample after having filtered out statistically insignificant ones using techniques such as Kernel Density Estimation (KDE) or Distance To a Measure (DTM) as suggested in Chazal et al. (2011); Wasserman (2016). Nevertheless, given the limited number of nearest neighbours in the vicinity of the query it will likely be challenging to derive statistically meaningful signals from higher order homologies.
5.2 Geometry intuition behind homology signatures
Unambiguous or clear samples appear to distribute narrowly around $W_{1}(H_{0})=0.6$. Given the fact that a single topic cluster is involved in this case, it is possible to envision a simple simulation (detailed in appendix B) in order to link the norm distribution in terms of neighbourhood orientation anisotropy. The observed behaviour from the data and the simulation is an anisotropy of neighbourhood, corresponding to an anisotropy factor $\alpha=15$ (refer to appendix B) for the definition). One topological consequence is that the query is located outside the convex hull created by the (k-1)-dimensional simplex generated by the k neighbours. It appears to be a geometric signature of queries with higher clarity nevertheless not sufficient to fully attribute it. In the context of unknown topics clusters multiplicity in the neighbourhood, our simulation shows that the multiplicity of clusters distributed across orthogonal dimensions and their corresponding effective anisotropies vary significantly for a single $W_{1}(H_{0})$ value. Therefore, if in the regions $W_{1}(H_{0})\in[0.2,0.4]\cup[0.9,1.0]$ this single feature is an acceptable signal, it gets more complex between those two intervals and $H_{1}$ signature appears then to be needed to discriminate clear from ambiguous queries at various degrees. With similar apparent $H_{0}$ feature value, indeed long-lived loops, characterised by $LT_{max}(H_{1})$, appear to be more probable in the case of ambiguous queries in our experiment.
The presence such long lived loops could approximate discontinuities in the manifold underlying the neighbourhood in the case of ambiguous queries. It is broadly accepted that Language Models encode embeddings on a low dimensional sub-space. Manifold smoothing strategies for SBERT as well as recent successes of highly compressed models (Lee et al. (2024),Kusupati et al. (2024)) add empirical evidence to that conjecture. The apparent high degree of anisotropy of the projected neighbourhood of a query onto a sphere $S^{n}$ seems in agreement with a set of neighbours contained in a much lower dimensional sphere $S^{m}$ with $m\ll n$, compatible with a semantic manifold vision. Under this assumption, the presence of persistent loops could corresponds to punctures in the manifold. As a consequence, it would constitute semantic discontinuities between the query neighbours, thus reinforcing the idea that the query lies in a manifold singularity. One can note rule out either that the embedding space $\mathbb{R}^{n}$ of the tested model here contain more than a single low dimensional semantic manifolds and that the query lies at their intersection.
5.3 Perspectives: the BLOWFISH toolbox
We have shown empirical evidence for ambiguity signatures in query-chunk homology derived features. Supported by the experimental results, we believe that those features and others derived from $H_{0}$ and $H_{1}$ could be used to improve semantic search, particularly as they are embedding model architecture and training data agnostic (to the extent of the models we could test). We propose here a methodology for detecting ambiguity in sentences using homology signatures and kernel density estimation (KDE) under the name of BLOWFISH toolbox. This method allows us to express ambiguity as a probability for effective ambiguity/clarity detection. It builds on evidence from controlled experiments and incorporates several components for enhancing the detection process.
The BLOWFISH toolbox
Components
1. Homology Signatures
Identify homological signatures $H_{0}$ and $H_{1}$ for a specified scaling factor $\varepsilon$: $f_{homology}(\varepsilon)$
2. Ambiguity Score construction
Construct an ambiguity score based on homology features: $P(A|f_{homology}(\varepsilon),\rho_{queries}(\varepsilon),\theta_{embedding})\equiv\text{KDE}(f_{homology}(\varepsilon),\rho_{queries}(\varepsilon),\theta_{embedding})$ where $P(A|.)$ is the probability that a query is ambiguous, $\rho_{queries}(\varepsilon)$ is the density of neighbours for a particular scaling factor $\varepsilon$ (defined in appendix A.4) and $\theta_{embedding}$ representing the weights of the embedding model.
3. Model Independence
The ambiguity score is empirically independent of the embedding model of the datapoints it is applied to: $P(A|f_{homology}(\varepsilon),\rho_{queries}(\varepsilon),\theta_{embedding})\rightarrow P(A|f_{homology}(\varepsilon),\rho_{queries}(\varepsilon))$
Notably, the ambiguity score is independent of the embedding model used. This independence enables the use of smaller, cost-effective embedding models for fitting the KDE, which can then be applied to larger models, making the process efficient and scalable. It is also important to emphasize that whereas we ruled out the use of synthetic queries-answers datasets for this work which focuses on identifying signals, we believe that this method could be leveraged to build ambiguity score models with robust statistics.

### v2

@@ 3.3 Understanding Wasserstein Distance in Topological Data Analysis
The Wasserstein distance, also known as the earth mover’s distance, quantifies the ”work” required to transform one distribution into another. In topological data analysis, this measure is particularly useful when applied to persistence diagrams. The diagonal of a persistence diagram ideally represents features (often considered as noise) that appear and disappear at the same scale, thus serving as a baseline for comparison.
@@
where $N$ is the number of points in $H_{0}$, and $\gamma_{\perp}(y)$ is the y-coordinate of the orthogonal projection of each point onto the diagram’s diagonal. This calculation measures how far each point deviates from the baseline (diagonal), it implicitly weighs the features by their granularity. Each distance moved in the calculation represents a weighted adjustment of the data’s granularity, reflecting varying scales of data clustering and feature resolution.
@@ 3.4 Maximum loop lifetime
3.5 ”Ab-initio” simulation
We have created ”ab-initio” simulations of ambiguity as defined in the introduction on the sole linear representation hypothesis experiment. In this framework, random topics are represented as a vocabulary of normalized and orthogonal vectors in an arbitrary embedding space and datapoints are a linear combinations of those topics. For simplicity We assume uniform distribution of the topics that compose a datapoint.
3.5.1 Simulation design
The topics vocabulary is composed of $N$ different topics and the arbitrary embedding space has $D$ dimensions. We can therefore sample datapoints containing $n$ features and build pairs of queries and corpus datapoints. We designed 3 scenario that model the definitions of polysemy and multi-factuality and derived the persistent homologies results in various conditions for the hyperparameters ($D,N,n_{\text{parent}},\varepsilon,\sigma_{\text{noise}}$):
Scenario 1 hypothesis, pure multi-factual regime: Parent queries (topics set $S^{\text{parent}}_{i}$) retrieving children (topics sets $\{S^{\text{child}}_{i,j}:S^{\text{child}}_{i,j}\subset S^{\text{parent}}_{i}\}$ ) should show strong connectivity as children inherit parent topics, as children embeddings lie in a subspace spanned by the parent’s topics.
Scenario 2 hypothesis, pure polysemic regime: Child queries (topics sets $\{S^{\text{child}}_{i,j}:S^{\text{child}}_{i,j}\subset S^{\text{parent}}_{i}\}$) retrieving parents ($\{S^{\text{parent}}_{i}\}$) may show weaker connectivity as parent embeddings span additional topic dimensions beyond those present in the child.
Scenario 3 hypothesis, locally multi-factual: Mixing quasi-orthogonal children (from different parent lineages) with non-orthogonal children (from same parent) should produce intermediate scenario where topological structure varies significantly from local to global neighborhood.
Each corpus sample is composed of 50 datapoints. A detailed explanation of the algorithm can be found in appendix B.
3.6 Experiment Methodology
The challenge of evidencing consistent homology signal in an real dataset is the lack of prior information on the topics distribution in the corpus. We adopted a proxy strategy consisting of chunking queries and answers with different chunk sizes to approximate the number of topics in the query and corpus. Again, we motivate this choice with the empirical evidence that text sequences are constructed from distributions over topics where a sequence contains a set of topics and a subsequence contains a subset of those topics. It has the advantage of correlating the amount of topics in queries and answers and therefore limiting potential confounding factors in the observed topological signal. Synthetic queries generation does not guarantee such a degree of control to evidence the meaningful signals which explain why we have not chosen this popular dataset generation.
We conducted our experiments using the a set of Nobel Prize Physics lectures from 1901 to 2024, available as open-access transcripts on the Nobel Prize website. The corpus contains 14 lectures and about 250000 tokens in total. Each lecture was segmented into contiguous, non-overlapping chunks at two granularity : 250 (“fine”) and 750 (“coarse”). This produced approximately 919 and 320 chunks respectively (Table 1).
Dataset Chunk size (tokens) # chunks
$\mathcal{D}_{750}$ 750 320
$\mathcal{D}_{250}$ 250 919
Table 1: Nobel-lectures corpus: chunk granularities and counts
We conducted semantic searches using cosine similarity metric for each model and dataset, and collected the top k = 50 retrieved answers for each query. We encoded the content of these datasets using four different embedding models, as detailed in Table 2, then stored each resulting vector embedding in separate FAISS (Douze et al. (2024)) vector store indexes. These models were selected due to the diversity of corpora used in their pre-training and the number of embedding dimensions. None of these models were fine-tuned on our datasets for the experiment.
Embedding model Vector dimension
text-embedding-005 768
text-embedding-v3-small 1536
text-embedding-v3-large 3072
tasksource/ModernBERT-base-embed 768
Table 2: Embedding models used in the experiments
The generic experimental design shown is the following:
  1. 1.
    Chunk formation Divide the dataset $\mathcal{D}$ into chunks of size (number of tokens) $T\in\mathcal{I}=\{250,\,750\}$. This produces the two test chunk sets $\{\mathcal{C}_{T}\}_{T\in\mathcal{I}}$ listed in Table 1 with chunking consistency that requests $\bigcup_{c\in\mathcal{C}_{250}}c=\bigcup_{c\in\mathcal{C}_{750}}c=\mathcal{D}$.
  2. 2.
    Setup Use one set as the corpus $\mathcal{C}_{T}=\mathcal{C}$ and sample queries from the other set $\mathcal{Q}\sim\mathcal{C}_{T^{\prime}}$ as the query set $\mathcal{Q}$. Generate embeddings for both $\mathcal{Q}$ and $\mathcal{C}$, index $\mathcal{C}$
  3. 3.
    Retrieval Perform semantic search on corpus chunks for each query and keep the top-$k=50$ retrieved corpus chunks.
  4. 4.
    Evaluation Compute persistent homology features and test their ability to separate ambiguous from unambiguous queries.
@@ 4 Results
4.1 Simulation
Our ”ab-initio” simulation shows a clear separation between scenarios 1 and 2 for both homology metrics $W_{1}(H_{0})$ and $LT_{max}(H_{1})$ (figure 1 - Top) with an average separation of $\Delta W_{1}(H_{0})$0.2 and $\Delta LT_{max}(H_{1})=0.05$. Their value as a function of the scaling of the neighborhood $\varepsilon$ stay separated for both metrics 1 - Bottom). Scenario 1 metrics stabilize around $\varepsilon=1.0$ which corresponds to the scale where all 50 corpus datapoints are captured in the neighborhood. Scenario 2 metrics stabilize further away, around $\varepsilon=2.0$. It is consistent with the fact that scenario 1 only consists in local neighborhood (all corpus datapoints have a subset of topics present in the query) while in scenario 2, corpus datapoints embeddings are unbounded by the query topics as they either consist in a super-set of those topics or they are uncorrelated.
Figure 1: Top Left) Simulated $W_{1}(H_{0})$ for scenario 1 (red) and scenario 2 (blue) with neighborhood scale $\varepsilon=0.4$. Top Right) Simulated $LT_{max}(H_{1})$ for scenario 1 (red) and scenario 2 (blue) with neighborhood scale $\varepsilon=0.4$. Bottom Left) Simulated $W_{1}(H_{0})$ for scenario 1 (red) and scenario 2 (blue) as a function of the relative neighborhood scaling factor $\varepsilon$. Bottom Right) Simulated $LT_{max}(H_{1})$ for scenario 1 (red) and scenario 2 (blue) as a function of the relative neighborhood scaling factor $\varepsilon$. Simulation conditions: ($D=256,N=64,n_{\text{parent}}=32,\varepsilon=0.4,\sigma_{\text{noise}}=0.1$)
While simulation scenarios 1 and 2 represent the two extrema of conditions, scenario 3 is a mixture of strongly and loosely correlated datapoints that should be a more faithful approximation of real world case. We can see that the $W_{1}(H_{0})$ signal latter evolves more smoothly with $\varepsilon$ than scenario 1 (figure 2-Left) but is still separated from scenario 2 in absolute value, still allowing to determine the ambiguity regime. Similarly, absolute value for $LT_{max}(H_{1})$ decreases and does not show as sharp increase as a function of $\varepsilon$ (figure 2-Right). Additional results in different simulation conditions (including embedding dimensions and projections) are presented in appendix B and show robustness of the metrics.
Figure 2: Left) Simulated $W_{1}(H_{0})$ for scenario 1 (plain) and scenario 3 (dashed) as a function of the relative neighborhood scaling factor $\varepsilon$. Right) Simulated $LT_{max}(H_{1})$ for scenario 1 (plain) and scenario 2 (dashed) as a function of the relative neighborhood scaling factor $\varepsilon$. Simulation conditions: ($D=256,N=64,n_{\text{parent}}=32,\varepsilon=0.4,\sigma_{\text{noise}}=0.2$)
Therefore, under the joint assumptions that sequences set of topics sampled from a topic distribution (Grootendorst (2022)) and the linear representation hypothesis to encode those topics, we demonstrate that our formal definitions of ambiguity could be translated and quantified in a metric set using persistent homologies $W_{1}(H_{0})$ and $LT_{max}(H_{1})$. In order to validate those findings in practical applications we need to evidence such signals in real world dataset.
4.2 Experiment
In our experiments, the set $\mathcal{C}_{250}$ is initially selected as the corpus $\mathcal{C}$, with chunks from $\mathcal{C}_{750}$ used as the query set $\mathcal{Q}$. This configuration allows for the inference of query-related topics, because at least one “fine” chunk in $\mathcal{C}_{250}$ shares token sub-sequences with every query chunk drawn from $\mathcal{C}_{750}$. this setup corresponds to the multi-factual query case (scenario 2 in our simulation).
The experimental setup is reversed by selecting $\mathcal{C}_{750}$ as the corpus and $\mathcal{C}_{250}$ as the queries. This corresponds to the polysemic regime (scenario 1) where the queries are generally composed of less topics than the corpus chunks.
@@ 4.2 Experiment
To maintain the integrity of our experiments, we stipulate that the top retrieved chunk, denoted as $c_{i}=[w^{i}_{1},\ldots,w^{i}_{m}]$, must fully contain or be contained by the query chunk in both experimental configurations $(\mathcal{Q};\mathcal{C})$: $(\mathcal{C}_{750};\mathcal{C}_{250})$ and $(\mathcal{C}_{250};\mathcal{C}_{750})$. This ensures that the semantic coherence between the chunks is adequately maintained, allowing for accurate assessment of query ambiguity and topic distribution within the chunks.
@@
Figure 3: Left) $W_{1}(H_{0})$ KDE plot distributions for pairs of queries-corpus $(\mathcal{Q};\mathcal{C})$ chunk sets ($\mathcal{C}_{250}$$\mathcal{C}_{750}$) and ($\mathcal{C}_{750}$$\mathcal{C}_{250}$) in blue and red respectively for neighborhood scale $\varepsilon=0.4$. Right) Max $H_{1}$ loop lifetime KDE plot distributions for ($\mathcal{C}_{250}$$\mathcal{C}_{750}$) and ($\mathcal{C}_{750}$$\mathcal{C}_{250}$) in blue and red respectively with neighborhood scale $\varepsilon=0.4$.
@@
Figure 4: Left) Mean $W_{1}(H_{0})$ (plain lines) and 25-75% inter-quantile range (dashed lines) for ($\mathcal{C}_{250}$$\mathcal{C}_{750}$) and ($\mathcal{C}_{750}$$\mathcal{C}_{250}$) in blue and red respectively as a function of the relative neighborhood scaling factor $\varepsilon$. Right) equivalent plot for $LT_{max}(H_{1})$. The presented data is for ModernBERT-base-embed model.
In the experimental pairing of $\mathcal{Q}=\mathcal{C}_{250}$ and $\mathcal{C}=\mathcal{C}_{750}$, where the predefined conditions of polysemy are met, the $H_{0}$ signal is in the region $W_{1}(H_{0})=0.2-0.5$ (figure 3-Left). In contrast, the pair $\mathcal{Q}=\mathcal{C}_{750}$ and $\mathcal{C}=\mathcal{C}_{250}$ exhibits a shifted distribution to a higher value of $W_{1}(H_{0})$ by a consistent $\approx+0.2$ across all models tested. This shift is a partial indication of a more structured neighborhood. We cannot observe any obvious pattern related as a function of the embedding size of the models and all results appear in agreement with the simulation. Similarly, as the neighborhood scaling factor $\varepsilon$ increases, both distributions shift to lower values while maintaining the same mean distance (figure 4-Left) as in figure 1-Top.
@@
Same applies to $H_{1}$ which capture the lifetime of loops existing in the data points cloud. It exhibits more persistent elements (figure 3-Right) as expected from our model. As the neighborhood evolves from local to global (by varying $\varepsilon$), the ambiguous queries also start to exhibit persistent loops (figure 4-Right). Those loops can be interpreted as discontinuities that mark separation between different semantic clusters, reinforcing the picture of a pinched manifold that reflects a complex and discontinuous local embedding space surrounding the query chunks. the scaling is notably increased in the mean lifetime of unambiguous queries loops.
5 Discussion
5.1 Homology encodes ambiguity
We showed that persistent homology metrics significantly and consistently vary as a function of the query-corpus pairs conditions the query both in empirical setting on a real dataset with diverse range of embedding models as well as in an ”ab-initio” simulation where data and encoding functions are abstracted and only relying on the principle of linear representation of features/topics.
This behavior is observed across all 4 different embedding models both for $H_{0}$ and $H_{1}$ and in agreement with our simulation demonstrating that topological signatures of ambiguity are invariant with embedding model architecture variation and training corpus. The agreement on $W_{1}(H_{0})$ between simulation and ModernBERT embeddings is remarkable. It is important though to discuss the shift upward for the 3 other models. They have in common of having been trained with the Matryoshka Representation Learning loss as a regularization function (Kusupati et al. (2024), Lee et al. (2024)). It optimizes the representation of a sequence of text in subspaces of decreasing dimensions all encapsulated in the same vector and that diverges strongly from the construction of datapoints in our simulation. ModernBERT, on the other hand is exempt from regularization. Despite this shift, polysemic and multi-factual regimes are well separated in all models.
In a practical case where the specifics of the embedding model are unknown, one could perform a calibration step of ambiguous vs unambiguous persistent homology values using the same methodology inspired by our real data experiment:
  • Identify clusters of documents.
  • Sample documents per cluster to create multi-factual queries and embed them
@@ 5.1 Homology encodes ambiguity
Those results allows us to state that persistent homologies $H_{0}$ and $H_{1}$ are only functions of the polysemy of the queries and therefore encode the relative ambiguity of a query with a specific neighborhood. Nevertheless, given the limited number of nearest neighbors in the vicinity of the query it will likely be challenging to derive statistically meaningful signals from higher order homologies.
Our findings underscore $H_{0}$ and $H_{1}$ homology signatures captured by $W_{1}(H_{0})$ and $LT_{max}(H_{1})$ as powerful tools to differentiate the two sets of query-corpus we explored. The design of our experiment aligns with our formal definition of sentence ambiguity/clarity.
5.2 Retrieval recall optimisation
As discussed above, absolute values of our homology metrics encode ambiguity by defining which regime the search query is in. If the query is multi-factual it is also important to determine the optimal top k retrieval value. It is easy to identify that $W_{1}(H_{0})$ has the same global trend in the experiment and in the simulation as the neighborhood expand but the experimental result show non monotonous behavior of $dW_{1}(H_{0},\varepsilon)/d\varepsilon$ resulting in the existence of different slopes. We attribute this discrepancy to the ”purity” of the corpus created in the simulation versus the real world case: that is to say all corpus datapoints are either strongly correlated to the query (scenario 1) or loosely correlated (scenario 2) so that the distribution of degrees of correlation between the retrieved chunks and the query is bi-modal. On the contrary, in the experiment the distribution of degrees of correlation between the retrieved chunks and the query is likely more complex. Scenario 3 simulation illustrate that neighborhood heterogeneity is encoded in the homology metrics evolution as a function of $\varepsilon$. Therefore, what we observe here is could be used to perform hierarchical clustering of the retrieved chunks as a function of epsilon akin to HDBSCAN (McInnes et al. (2017)) and to optimize recall by dynamically updating $k$ the number of retrieved chunks to return.
Furthermore we state that the question of recall optimization is only relevant in the case where the query is multi-factual. Indeed, a polysemic query by definition lacks of information and recalling loosely correlated datapoints would only inject noise as an output of the retrieval process and in RAG systems it can lead to generative model misrepresentation of the context and hallucination. For such a query, the best approach would be an exception handling asking for a reformulation with more context.
So far we have presented multi-factual queries simulation results for (scenario 1) and highlighted the similarities with the real world experimental results. We exhibited differences between the two settings, especially for $LT_{max}(H_{1})$ and attributed it to ”purity” of the corpus in the simulation.